OpenStudy (anonymous):

Will give medal! How can you tell when an expression is undefined? Some examples: sec (-pi/2) csc (3*pi) cot (7*pi/2) csc (-3*pi/2) cot (5*i/3)

OpenStudy (jdoe0001):

so.... when is a fraction undefined?

OpenStudy (anonymous):

when zero is the denominator, but I don't see how you can tell that in this situation. I've tried typing them all into a calculator, but that doesn't help.

OpenStudy (anonymous):

When the denominator is 0.

OpenStudy (anonymous):

Remember that pi=180. You need to simplify.

OpenStudy (anonymous):

I thought pi was equal to 3.14 rounded? Sorry for the late response. I was looking at another question.

OpenStudy (jdoe0001):

hmm yes when the denominator is 0 so

OpenStudy (anonymous):

So pi = 180 when the denominator equals 0?

OpenStudy (jdoe0001):

one sec

OpenStudy (anonymous):

okay

OpenStudy (jdoe0001):

\(\bf sec\left(-\frac{\pi}{2}\right)\implies \cfrac{1}{{\color{brown}{ cos}}\left(-\frac{\pi}{2}\right)}\implies \cfrac{1}{{\color{brown}{ cos}}\left(\frac{\pi}{2}\right)} \\ \quad \\ csc\left(\frac{3\pi}{2}\right)\implies \cfrac{1}{{\color{brown}{ sin}}\left(\frac{3\pi}{2}\right)} \\ \quad \\ cot\left(\frac{7\pi}{2}\right)\implies \cfrac{cos\left(\frac{7\pi}{2}\right)}{{\color{brown}{ sin}}\left(\frac{7\pi}{2}\right)} \\ \quad \\ csc\left(-\frac{3\pi}{2}\right)\implies \cfrac{1}{{\color{brown}{ sin}}\left(-\frac{3\pi}{2}\right)}\implies \cfrac{1}{{\color{brown}{ -sin}}\left(\frac{3\pi}{2}\right)}\) so.. if you find the value of the function in the denominator.. and if it zeros out....then the fraction will be undefined, and so will the trig function

OpenStudy (jdoe0001):

and the last one, the cotangent one, is the same as that one above just change it to cosine and sine based and check if at that angle the denominator zeros out

OpenStudy (anonymous):

So cot (3*pi/2) would be undefined since it cancels out?

OpenStudy (jdoe0001):

you mean csc?

OpenStudy (anonymous):

yeah sorry csc

OpenStudy (jdoe0001):

\(\bf csc\left(\frac{3\pi}{2}\right)\implies \cfrac{1}{{\color{brown}{ sin}}\left(\frac{3\pi}{2}\right)}\implies \cfrac{1}{-\cancel{ 1 }}\to -1\)

OpenStudy (jdoe0001):

anyhow they cancel out.. but \(\bf \cfrac{\cancel{ 1 }}{-\cancel{ 1 }}\to -1 \)

OpenStudy (jdoe0001):

one of the ones that's simple will be say the 1st one \(\bf sec\left(-\frac{\pi}{2}\right)\implies \cfrac{1}{{\color{brown}{ cos}}\left(-\frac{\pi}{2}\right)}\implies \cfrac{1}{{\color{brown}{ cos}}\left(\frac{\pi}{2}\right)}\to \cfrac{1}{0} \)

OpenStudy (anonymous):

When I typed them into the calculator cos(pi*2) = 0 cos(7*pi/2)/sine(7*pi/2) = 0

OpenStudy (anonymous):

and based off what you said earlier that means that cot (5*i/3) = 0 as well correct?

OpenStudy (anonymous):

My assignment says that: sec(-pi/2) and csc(3*pi) are the only undefined expressions. Could you tell me where I went wrong?

OpenStudy (jdoe0001):

well... \(\bf cot\left(\frac{5\pi}{3}\right)\implies \cfrac{cos\left(\frac{5\pi}{3}\right)}{sin\left(\frac{5\pi}{3}\right)}\to \cfrac{\frac{1}{2}}{-\frac{\sqrt{3}}{2}}\to \cfrac{1}{\cancel{ 2 }}\cdot -\cfrac{\cancel{ 2 }}{\sqrt{3}}\)

OpenStudy (anonymous):

Ah I see now. but how does cot (7*pi/2) not equal zero?

OpenStudy (jdoe0001):

ok well keep in mind that \(\bf \cfrac{7}{2}\to \cfrac{6}{2}+\cfrac{1}{2}\to 3+\cfrac{1}{2}\) so one can say that \(\bf \cfrac{7\pi}{2}\to \cfrac{6\pi}{2}+\cfrac{\pi}{2}\to 3\pi+\cfrac{\pi}{2}\) so. where's that angle anyway? well is 3 go-round or full cycles and then \(+\cfrac{\pi}{2}\) so